# SpatialActor: Exploring Disentangled Spatial Representations for Robust Robotic Manipulation

### v1

@@ Abstract
Robotic manipulation requires precise spatial understanding to interact with objects in the real world. Point-based methods suffer from sparse sampling, leading to the loss of fine-grained semantics. Image-based methods typically feed RGB and depth into 2D backbones pre-trained on 3D auxiliary tasks, but their entangled semantics and geometry are sensitive to inherent depth noise in real-world that disrupts semantic understanding. Moreover, these methods focus on high-level geometry while overlooking low-level spatial cues essential for precise interaction. We propose SpatialActor, a disentangled framework for robust robotic manipulation that explicitly decouples semantics and geometry. The Semantic-guided Geometric Module adaptively fuses two complementary geometry from noisy depth and semantic-guided expert priors. Also, a Spatial Transformer leverages low-level spatial cues for accurate 2D-3D mapping and enables interaction among spatial features. We evaluate SpatialActor on multiple simulation and real-world scenarios across 50+ tasks. It achieves state-of-the-art performance with 87.4% on RLBench and improves by 13.9% to 19.4% under varying noisy conditions, showing strong robustness. Moreover, it significantly enhances few-shot generalization to new tasks and maintains robustness under various spatial perturbations. Project Page: https://shihao1895.github.io/SpatialActor
@@ 2.2 Vision Foundation Models for Robotics
Vision foundation models have significantly enhanced robotic perception by incorporating semantic and geometric priors. Visual and multimodal models (Radford et al. 2021; Li et al. 2022; Feng et al. 2023; Liu et al. 2024b; Wang et al. 2025b) leverage diverse datasets to learn strong semantic priors that improve visual understanding, which benefits downstream robotic tasks. However, they primarily focus on the 2D domain and lack spatial understanding capabilities. 3D vision models (Zhu et al. 2024; Zheng et al. 2024; Qian et al. 2022; Zheng et al. 2025; Kang et al. 2024; Zhang et al. 2025) integrate semantic information with explicit spatial structures to facilitate effective geometric perception. However, the acquisition and annotation of 3D data are inherently expensive and labor-intensive, which restricts scalability and limits their application in real-world scenarios. Depth estimation experts (Yang et al. 2024, 2025; Bhat et al. 2023; Wang et al. 2025a) leverage large-scale pretraining on diverse datasets to translate semantics in images into corresponding geometric structures, robustly inferring geometric information even under challenging conditions such as sensor noise and occlusions. In this paper, we leverage the strong semantic alignment of vision models together with robust geometric priors from depth estimation experts.
@@ 4.4 Spatial Perturbations on ColosseumBench
Experimental Setup.
We evaluate the robust capability of SpatialActor on the Colosseum benchmark (Pumacay et al. 2024), which is designed to assess robot manipulation policies under environmental changes. We evaluate performance on 20 benchmark tasks under both baseline (no perturbation) and three spatial perturbation conditions. The first, manipulation object size (MO-Size), scales the object the robot directly interacts with, simulating dimensional variations. The second, receiver object size (RO-Size), alters the size of an indirectly used object, such as a container or support surface. The third condition introduces camera pose perturbations by randomly adjusting the camera’s position and orientation, mimicking changes in the observation viewpoint. Together, these perturbations capture spatial environmental variations that can affect robot performance.
@@ 4.5 Ablation Study
Our ablation study on 18 tasks (Tab. 5) shows that decoupling semantics and geometry improves performance in both noise-free and heavy-noise settings, increasing success rates from 81.4% to 85.1% and from 57.0% to 68.7%, respectively. Introducing the Semantic-guided Geometry Module (SGM) further boosts performance, especially under heavy noise, where performance rises to 73.9%. Finally, the Spatial Transformer (SPT), which provides precise low-level spatial cues, brings the success rates to 87.4% and 76.4% in noise-free and noisy conditions, respectively. These results highlight the importance of each proposed component in improving both accuracy and robustness.
4.6 Real-World Evaluation
Setup.
In real-world experiments, we use a WidowX single-arm robot equipped with an Intel RealSense D435i RGB-D camera. The camera is statically mounted to capture a front view of the workspace. We perform both intrinsic and extrinsic calibration between the camera and the robot to accurately transform the observed point clouds into the robot’s base coordinate system, ensuring precise manipulation. The system is integrated using a ROS package (Coleman et al. 2014). Images are originally captured at a resolution of $1280\times 720$ and are downsampled to $128\times 128$.
@@ 4.6 Real-World Evaluation
Table 4: Performance Under Spatial Perturbations. We report average success rates on 20 ColosseumBench tasks under four conditions: No-Vars, manipulation object size (MO-Size), receiver object size (RO-Size), and camera pose (Cam Pose). SpatialActor achieves higher performance.
@@
Task #vari. RVT-2 SpatialActor (Ours)
@@ 5 Conclusion
In this work, we present SpatialActor, a framework for robust spatial representation in robotic manipulation that addresses the challenges of precise spatial understanding, sensor noise, and effective interaction. SpatialActor disentangles semantic and geometric information, with the geometric branch divided into high-level and low-level components: SGM adaptively fuses semantic-guided geometric priors with raw depth features for robust high-level geometry, while SPT captures low-level spatial cues through position-aware interactions. Extensive experiments across 50+ simulated and real-world tasks demonstrate that SpatialActor achieves higher success rates and strong robustness under diverse conditions. These results highlight the importance of disentangled spatial representations for developing more robust and generalizable robotic systems.
@@ Appendix H Simulation Tasks
We follow the evaluation protocol of PerAct and benchmark SpatialActor on the same 18 RLBench tasks shown in Figure 10. In total, these tasks cover 249 randomized scene configurations varying in object color, count, placement, shape, size, or category. Table 11 provides each task’s language instruction template, the number of variations, and the type of variation.

### v2

@@ Abstract
Robotic manipulation requires precise spatial understanding to interact with objects in the real world. Point-based methods suffer from sparse sampling, leading to the loss of fine-grained semantics. Image-based methods typically feed RGB and depth into 2D backbones pre-trained on 3D auxiliary tasks, but their entangled semantics and geometry are sensitive to inherent depth noise in real-world that disrupts semantic understanding. Moreover, these methods focus on high-level geometry while overlooking low-level spatial cues essential for precise interaction. We propose SpatialActor, a disentangled framework for robust robotic manipulation that explicitly decouples semantics and geometry. The Semantic-guided Geometric Module adaptively fuses two complementary geometry from noisy depth and semantic-guided expert priors. Also, a Spatial Transformer leverages low-level spatial cues for accurate 2D-3D mapping and enables interaction among spatial features. We evaluate SpatialActor on multiple simulation and real-world scenarios across 50+ tasks. It achieves state-of-the-art performance with 87.4% on RLBench and improves by 13.9% to 19.4% under varying noisy conditions, showing strong robustness. Moreover, it significantly enhances few-shot generalization to new tasks and maintains robustness under various spatial perturbations.
Project Pagehttps://shihao1895.github.io/SpatialActor
Codehttps://github.com/shihao1895/SpatialActor
@@ 2.2 Vision Foundation Models for Robotics
Vision foundation models have significantly enhanced robotic perception by incorporating semantic and geometric priors. Visual and multimodal models (Radford et al. 2021; Li et al. 2022; Feng et al. 2023; Liu et al. 2024b; Wang et al. 2025b; Wu et al. 2025) leverage diverse datasets to learn strong semantic priors that improve visual understanding, which benefits downstream robotic tasks. However, they focus on the 2D domain and lack spatial understanding capabilities. 3D vision models (Zhu et al. 2024; Zheng et al. 2024; Qian et al. 2022; Zheng et al. 2025; Kang et al. 2024; Zhang et al. 2025) integrate semantic information with explicit spatial structures to facilitate effective geometric perception. However, the acquisition and annotation of 3D data are inherently expensive and labor-intensive, which restricts scalability and limits their application in real-world scenarios. Depth estimation experts (Yang et al. 2024, 2025; Bhat et al. 2023; Wang et al. 2025a) leverage large-scale pretraining on diverse datasets to translate semantics in images into corresponding geometric structures, robustly inferring geometric information even under challenging conditions such as sensor noise and occlusions. In this paper, we leverage the strong semantic alignment of vision models together with robust geometric priors from depth estimation experts.
@@ 4 Experiments
Models Put in Cupboard Put in Drawer Put in Safe Screw Bulb Slide Block Sort Shape Stack Blocks Stack Cups Sweep Dustpan Turn Tap
@@ 4.4 Spatial Perturbations on ColosseumBench
Setup.
We evaluate the robustness on the Colosseum benchmark (Pumacay et al. 2024), which assesses manipulation policies under environmental variations. We report results on 20 tasks under the baseline setting (no perturbation) and three spatial perturbations: manipulation object size (MO-Size), which scales the object being manipulated; receiver object size (RO-Size), which scales an indirectly used object such as a container; and camera pose perturbation (Camera-Perturb), which randomly shifts the camera’s position and orientation to vary the observation viewpoint.
@@ 4.5 Ablation Study
Our ablation study on 18 tasks (Tab. 5) shows that decoupling semantics and geometry improves performance in both noise-free and heavy-noise settings, increasing success rates from 81.4% to 85.1% and from 57.0% to 68.7%, respectively. Introducing the Semantic-guided Geometry Module (SGM) further boosts performance, especially under heavy noise, where performance rises to 73.9%. Finally, the Spatial Transformer (SPT), which provides precise low-level spatial cues, brings the success rates to 87.4% and 76.4% in noise-free and noisy conditions, respectively.
@@
Table 4: Performance Under Spatial Perturbations. We report average success rates on 20 ColosseumBench tasks under 4 conditions: No-Vars, manipulation object size, receiver object size, and camera pose.
4.6 Real-World Evaluation
Setup.
In real-world experiments, we use a WidowX single-arm robot equipped with an Intel RealSense D435i RGB-D camera. The camera is statically mounted to capture a front view of the workspace. We perform both intrinsic and extrinsic calibration between the camera and the robot to accurately transform the observed point clouds into the robot’s base coordinate system. The system is integrated using a ROS package. Images are originally captured at a resolution of $1280\times 720$ and are downsampled to $128\times 128$.
@@ 4.6 Real-World Evaluation
Task #Var. RVT-2 SpatialActor
@@ 5 Conclusion
In this work, we present SpatialActor, a framework for robust spatial representation in robotic manipulation. SpatialActor disentangles semantic and geometric information, with the geometric branch comprising a high-level module (SGM), which fuses semantic-guided geometric priors with depth features, and a low-level module (SPT), which captures fine-grained spatial cues through position-aware interactions. Experiments across 50+ simulated and real-world tasks show that SpatialActor achieves higher success rates and strong robustness, underscoring the value of disentangled spatial representations for reliable manipulation.
@@ Appendix H Simulation Tasks
We follow the evaluation protocol of PerAct and benchmark SpatialActor on the same 18 RLBench tasks shown in Figure 10. In total, these tasks cover 249 randomized scene configurations varying in object color, count, placement, shape, size, or category. Table 11 provides each task’s language instruction template, the number of variations, and the type of variation. Furthermore, Figure 11 illustrates the simulation tasks in a few-shot setting.
@@ Appendix I Real-world Tasks
Figure 11: Few-Shot tasks. We adapt pre-trained model to 19 new tasks using only 10 demonstrations per task (1/10th of original data).